OpenStudy (anonymous):

A company purchases shipments of machine components and uses this acceptance sampling plan: Randomly select and test 30 components and accept the whole batch if there are fewer than 3 defectives. If a particular shipment of thousands of components actually has a 7% rate of defects, what is the probability that this whole shipment will be accepted

OpenStudy (anonymous):

please help EMERGENCY!!!!!!! TIMED EXAM

OpenStudy (paxpolaris):

so if you take a sample of 30 ... you want either 0, 1, or 2 defective pieces

OpenStudy (anonymous):

I DONT UNDERSTAND

OpenStudy (paxpolaris):

you have a huge shipment (assume infinite) with 7 percent defective pieces.

OpenStudy (paxpolaris):

you take a sample of 30 pieces.. if 3 or more pieces are defective you cannot accept it

OpenStudy (paxpolaris):

so when you are taking a sample of 30 pieces ... you only want to find less than 3 defectives ... that is either 0, 1, or 2.

OpenStudy (paxpolaris):

do you follow so far.

OpenStudy (paxpolaris):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hello??????

OpenStudy (paxpolaris):

the probability of getting 0 defectives: \[(1-0.07)^{30}=(0.93)^{30}\]

OpenStudy (paxpolaris):

probability of 1 defective \[.93^{29}\times .07^1\times(_{30}C_1)\]

OpenStudy (paxpolaris):

probability of 2 defective :\[.93^{28}\times.07^2\times(_{30}C_2)\]